# Nonlinear Fore(Back)casting and Innovation Filtering for Causal-Noncausal VAR Models

### v3

@@ 7 Concluding Remarks
This paper considers the oos nonlinear forecasting and backcasting in a mixed VAR model. It introduces the closed-form expression of the causal (past-dependent) predictive distribution for forecasting a mixed (S)VAR model. As a post-estimation inference method, the confidence set of prediction set is introduced.
A definition of causal (past-dependent) nonlinear innovations for mixed VAR models is also given. Since the causal nonlinear innovations are not uniquely defined, their identification is examined.
The proposed approach is applied to the analysis of the joint dynamic of a bivariate series of cryptocurrency exchange rates. The state space representation reveals that the noncausal state variable captures the explosive patterns and can be interpreted as a crypto market portfolio, and that the causal state variable corresponds to a risk neutral fund. We also examined the impulse response functions associated with a common bubble shock (CBS).
@@
Andrews, B, Davis, R. and J., Breidt (2006): ”Maximum Likelihood Estimation of All-Pass Time Series Models”, Journal of Multivariate Analysis, 97, 1638-1659.
Axler, S., Bourdon, P. and W. Ramey (2001): ”Harmonic Function Theory”, 2nd Ed., Graduate Texts in Mathematics, Springer.
Bec, F., Nielsen, H., and S., Saidi (2020): ”Mixed Causal-Noncausal Autoregression: Bimodality Issues in Estimation and Unit Root Testing”, Oxford Bulletin of Economics and Statistics, 82, 1413-1428.
Bernoth, K., Herwartz, H., and L. Triener (2024): ”The Impact of Global Risk and US Monetary Policy on US Dollar Exchange Rates with Excess Currency Returns”, DP 2037, DIW, Berlin.
@@
Comon, P. (1994): ”Independent Component Analysis: A New Concept?”, Signal Process., 36, 287-314.
Cordoni, F., and F., Corsi (2019): ”Identification of Singular and Noisy Structural VAR Models: The Collapsing ICA Approach”, D.P. University of Pisa.
@@
Gonzalves, S., Herrera, A., Kilian, L., and E., Pesavento (2022): ”When Do State-Dependent Local Projections Work?”, McGill U, working paper.
@@
Gourieroux, C., Monfort A., and J.P., Renne (2020): ”Identification and Estimation in Nonfundamental Structural VARMA Models”, Review of Economic Studies, 87, 1915-1953.
@@
Guay, A. (2021): ”Identification of Structural Vector Autoregressions Through Higher Unconditional Moments”, Journal of Econometrics, 225, 27-45.
Hall, M., and J., Jasiak (2024): ”Modelling Common Bubbles in Cryptocurrency Prices”, working paper at www.jjstats.com
@@
Keweloh, S. (2021): ”A Generalized Method of Moments Estimator for Structural Vector Autoregressions Based on Higher Moments”, Journal of Business and Economic Statistics, 39, 772-782.
Koop, G., Pesaran, H., and S., Potter (1996): ”Impulse Response Analysis in Nonlinear Multivariate Models”, Journal of Econometrics, 74, 119-147.
@@
Lanne, M., and P., Saikkonen (2008): ”Modelling Expectations with Noncausal Autoregressions”, MPRA paper 8411.
Lanne, M., and P., Saikkonen (2011a) : ”Noncausal Autoregressions for Economic Time Series”, Journal of Time Series Econometrics, 3, 1-39.
Lanne, M., and P., Saikkonen (2011b) : ”GMM Estimators with Non-Causal Instruments”, Oxford Bulletin of Economics and Statistics, 71, 581-591.
@@
Lutkepohl, H. (1990): ”Asymptotic Distribution of Impulse Response Functions and Forecast Error Variance Decomposition of Vector Autoregressive Models”, Review of Economics and Statistics, 72, 116-125.
@@
Potter, S. (2000): ”Nonlinear Impulse Response Functions”, Journal of Economic Dynamics and Control, 24, 1425-1426.
@@
Starck, V. (2023): ”Estimation of Independent Component Analysis Systems”, Working Paper, Brown University.
@@
Twumasi, C. and J. Twumasi (2022): ”Machine Learning Algorithms for Forecasting and Backcasting Blood Demand Data with Missing Values and Outliers: A Study of Tema General Hospital of Ghana”, International Journal of Forecasting, 38, 1258-1277.
@@
APPENDIX A.1
Proof of Proposition 1
The proof follows along the same lines as in Gourieroux, Jasiak (2016), (2017), with an additional attention given to the information sets and the use of Jacobian formulas on manifolds. The derivation is greatly simplified for special cases of VAR(1) processes with a multiplicative representation [Lanne, Saikkonen (2013)].
@@
The multiplicative causal-noncausal model is:
$\Phi(L)\Psi(L^{-1})Y_{t}=\varepsilon^{*}_{t},$
where both autoregressive polynomials have roots outside the unit circle and i.i.d. errors $\varepsilon^{*}_{t}$ [Lanne, Luoto, Saikkonen (2012), Lanne, Saikkonen (2013) and Nyberg, Saikkonen (2014)]. While for univariate time series the multiplicative representation is equivalent to the general AR(p) model, it is not the case in the multivariate framework (except for the VAR(1) process).
As pointed out in Davis, Song (2020), p. 247, this decomposition implies restrictions on the autoregressive coefficients $\Phi_{1},...,\Phi_{p}$ of the past-dependent representation.
Moreover, it is not compatible in general with the VAR specification (2.1). To illustrate this problem, let us consider the multiplicative bivariate model:
$\left(\begin{array}[]{cc}1-\phi L&0\\ 0&1\end{array}\right)\left(\begin{array}[]{cc}1&0\\ 0&1-\psi L^{-1}\end{array}\right)Y_{t}=\varepsilon_{t}^{*}.$
It follows that:
$Y_{1,t}-\phi Y_{1,t-1}$ $=$ $\varepsilon_{1,t}^{*},$
$Y_{2,t}-\psi Y_{2,t+1}$ $=$ $\varepsilon_{2,t}^{*},$
or equivalently, if $\Psi\neq 0$,:
$Y_{1,t}-\phi Y_{1,t-1}$ $=$ $\varepsilon_{1,t}^{*},$
$Y_{2,t}-\frac{1}{\psi}Y_{2,t-1}$ $=$ $-\frac{1}{\psi}\varepsilon_{2,t-1}^{*},$
$\iff Y_{t}=\left(\begin{array}[]{cc}\phi&0\\ 0&\frac{1}{\psi}\end{array}\right)Y_{t-1}+\varepsilon_{t},$
where $\varepsilon_{t}=\left(\begin{array}[]{c}\varepsilon_{1,t}\\ \varepsilon_{2,t}\end{array}\right)$ with $\varepsilon_{1,t}=\varepsilon_{1,t}^{*}$ and $\varepsilon_{2,t}=-\frac{1}{\psi}\varepsilon_{2,t-1}^{*}$.
We observe that, if $\varepsilon_{1,t}^{*},\varepsilon_{2,t}^{*}$ are contemporaneously correlated, then $\varepsilon_{1,t}$ and $\varepsilon_{2,t+1}$ are correlated too. Therefore the condition of i.i.d. errors in the VAR model (2.1) cannot be satisfied.
@@
with i.i.d. errors, if matrix $\Psi$ is not invertible.
APPENDIX A.3
Identification of Nonlinear Causal Innovations
a) Existence of nonlinear causal autoregressive representation
For ease of exposition, let us consider a bivariate Markov process. By analogy to the recursive causal approach for defining the shocks, we start from the first component.
i) Let $F_{1}[y_{1}|Y_{T-1}]$ denote the conditional c.d.f. of $Y_{1,T}$ given $Y_{T-1}$ and define:
$v_{1,T}=F_{1}[Y_{1,T}|Y_{T-1}],\;\forall T.$ (a.7)
Then, $v_{1,T}$ follows a uniform distribution $U_{[0,1]}$ for any $Y_{T-1}$. In particular, $v_{1,T}$ is independent of $Y_{T-1}$.
ii) Let $F_{2}[y_{2}|Y_{1,T},Y_{T-1}]$ denote the conditional c.d.f. of $Y_{2,T}$ given $Y_{1,T},Y_{T-1}$, and define:
$v_{2,T}=F_{2}[Y_{2,T}|Y_{1,T},Y_{T-1}],\;\forall T.$ (a.8)
It follows that $v_{2,T}$ follows a uniform distribution on [0,1], for any $Y_{1,T},Y_{T-1}$, or equivalently for any $v_{1,T},Y_{T-1}$. Therefore, $v_{2,T}$ is independent of $v_{1,T},Y_{T-1}$.
iii) By inverting equations (a.6)-(a.7), we obtain a nonlinear autoregressive representation: $Y_{T}=a(Y_{T-1},v_{T})$, where the $v_{T}$’s are i.i.d. such that $(v_{1,T}),(v_{2,T})$ are independent.
Alternatively, one can use the ordering: $Y_{2,T}$ followed by $Y_{1,T}$ given $Y_{2,T}$. More generally, for any invertible nonlinear transformation $Y_{T}^{*}=c(Y_{T})$, the above approach can be applied first to $Y_{1,T}^{*}$ and next to $Y_{2,T}^{*}$ conditional on $Y_{1,T}^{*}$.
Therefore any Markov process can be written as a nonlinear causal autoregressive process and the above discussion shows that this autoregressive representation is not unique.
b) Identification of the nonlinear causal autoregressive representation
It is equivalent to consider the identification of function $a$ or the identification of nonlinear innovations. Let us now describe in detail all the nonlinear causal innovation identification issues. First, we can assume that $v_{1,T},v_{2,T}$ are i.i.d and independent of one another with uniform distributions on [0,1]. We need to find out if there exists another pair of variables $w_{1,T},w_{2,T}$, which are independent and uniformly distributed such that:
$a(Y_{T-1},w_{T})=a(Y_{T-1},v_{T}),\;\;\forall\;Y_{T-1},$
or, equivalently, a pair of variables $w_{T}$ that satisfy a (nonlinear) one-to-one relationship with $v_{T}$. Let $w=b(v)$ denote this relationship. We have the following Lemma:
Lemma A.1:
Let us assume that $b$ is continuous, twice differentiable and that the Jacobian matrix $\partial b(v)/\partial v^{\prime}$ has distinct eigenvalues. Then, the components of $b$ are harmonic functions, that is:
$\frac{\partial^{2}b_{j}(v)}{\partial v_{1}^{2}}+\frac{\partial^{2}b_{j}(v)}{\partial v_{2}^{2}}=0,\;\;j=1,2.$
Proof:
i) We can apply the Jacobian formula to get the density of $w$ given the density of $v$. Since both joint densities are uniform, it follows that $|\det\frac{\partial b(v)}{\partial v^{\prime}}|=1$, $\forall v\in[0,1]^{2}$.
ii) Let us consider the eigenvalues $\lambda_{1}(v),\lambda_{2}(v)$ of the Jacobian matrix $\frac{\partial b(v)}{\partial v^{\prime}}$. The eigenvalues are continuous functions of this matrix, and therefore continuous functions of $v$ (whenever these eigenvalues are different). Then, two cases can be distinguished:
case 1: The eigenvalues are real.
case 2: The eigenvalues are complex conjugates.
iii) In case 1, we have $\lambda_{2}(v)=1/\lambda_{1}(v)$ ( or $-1/\lambda_{1}(v)$), where $\lambda_{1}(v)$ is less or equal to 1 in absolute value for any $v$, and then $\lambda_{2}(v)$ is larger than or equal to 1 in absolute value for any $v$. Since $b(v)\in[0,1]^{2}$ for any $v\in[0,1]^{2}$, it follows that $\lambda_{2}(v)$ cannot be explosive.
iv) Therefore case 2 of complex conjugate roots is the only relevant one. Let us consider the case $\det\frac{\partial b(v)}{\partial v^{\prime}}=1$, $\forall v\in[0,1]^{2}$ (the analysis of $\det\frac{\partial b(v)}{\partial v^{\prime}}=-1$ is similar). Then, the Jacobian matrix is a rotation matrix:

### v4

@@ 7 Concluding Remarks
This paper considers oos nonlinear forecasting and backcasting in a mixed VAR model. It introduces a closed-form expression of forward (resp. backward) predictive density for forecasting (resp. backcasting) from mixed VAR models. As a post-estimation inference method, we adjust the estimated prediction interval by a conditional ”backward” bootstrap and introduce the confidence set of the estimated prediction interval. A definition of causal (past-dependent) nonlinear innovations for mixed VAR models is also given. Since the causal nonlinear innovations are not uniquely defined, their identification is examined and the IRF analysis following a shock to the noncausal component is discussed.
For illustration, the proposed approach is applied to the analysis of the joint dynamics of a bivariate series of oil prices and real GDP growth rates. We find that the noncausal state variable captures the explosive patterns, including bubbles and spikes. We examine the IRFs and observe different effects of a shock to the noncausal component on- and off-bubble. The mixed VAR models offers a parsimonious representation of nonlinear dynamic processes and allow for advanced analysis of the dynamics during a bubble episode.
Funding: This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC).
@@
Beresteanu, A., Molchanov, I., and F. Molinari (2017): ”Partial Identification Using Random Set Theory”, Journal of Econometrics, 16, 17-32.
Blasques, F., Koopman, S., Mingoli, G. and S., Telg (2025): ”A Novel Test for the Presence of Local Explosive Dynamics”, Journal of Time Series Analysis, forthcoming.
@@
Cubadda, G., Hecq, A., and S., Telg (2019): ”Detecting Co-Movements in Non-Causal Time Series, Oxford Bulletin on Economics and Statistics, 697-715.
@@
Cubadda, G., Giancaterini, F., Hecq, A. and J. Jasiak (2024): ”Optimization of the Generalized Covariance Estimator in Noncausal Processes”, Statistics and Computing, 34, 127.
@@
De Truchis, G., Freis, S. and A., Thomas (2025): ”Forecasting Extreme Trajectories Using Semi-Norm Representations”, working paper Paris Dauphine University.
@@
Giancaterini, F., Hecq, A., Jasiak, J. and A. Manafi-Neyazi (2025): ”Regularized Generalized Covariance Estimator”, ArXiv 6390395.
@@
Gourieroux, C., and Q., Lee (2025): ”Identification and Impulse Response Functions for Nonlinear Dynamic Models”, ArXiv 2506.13531.
@@
Hall, M., and J., Jasiak (2024): ”Modelling Common Bubbles in Cryptocurrency Prices”, Economic Modelling, 139, 106782.
@@
Herrera, A., Lagalo, L., and T., Wada (2015) : ”Asymmetries in the Response of Economic Activity to Oil Price Increases and Decreases ?”, Journal of International Money and Finance, 50, 108-133.
@@
Kilian, L., and R., Vigfusson (2017) : ”The Role of US Oil Price Shocks in Causing US Recessions”, Journal of Money, Credit and Banking, 40, 1747-1776.
@@
Lanne, M., and P., Saikkonen (2011) : ”Noncausal Autoregressions for Economic Time Series”, Journal of Time Series Econometrics, 3, 1-39.
@@
Lof, M., and H., Nyberg (2017): ”Noncausality and the Commodity Currency Hypothesis”, Energy Economics, 65, 424-433.
Molchanov, I., and F., Molinari (2018): ”Random Sets in Econometrics”, Econometric Society Monographs, Cambridge University Press.
@@
Rosenblatt, M. (1952) : ”Remarks on Multivariate Transformations”, Annals of Mathematical Statistics, 23, 470-472.
@@
Sims, C. (1980): ”Macroeconomics and Reality”, Econometrica, 48, 1-48.
@@
Twumasi, C., and J., Twumasi (2022): ”Machine Learning Algorithms for Forecasting and Backcasting Blood Demand Data with Missing Values and Outliers: A Study of Tema General Hospital of Ghana”, International Journal of Forecasting, 38, 1258-1277.
@@
APPENDIX A.1: Proof of Proposition 1
The proof employs the definition of information sets given in Gourieroux and Jasiak (2016), (2017). We derive the predictive density by using the Bayes theorem to condition the variables on their past and employ Jacobian formulas on manifolds.
@@
A constrained multiplicative Mixed Autoregressive (MAR) representation of the mixed VAR model was proposed by Lanne, Saikkonen (2013):
$\Phi(L)\Psi(L^{-1})Y_{t}=\epsilon_{t}^{*},$ (a.7)
where the matrices of autoregressive polynomials $\Phi$ and $\Psi$ have both roots outside the unit circle, such that $\Psi(0)=\Phi(0)=Id$ and $\epsilon_{t}^{*}$ is an i.i.d. sequence of errors. The multiplicative representation exists for the univariate mixed processes [Breidt et al. (1991), Lanne and Saikkonen (2008)]. For multivariate processes, Davis and Song (2020), Swensen (2022), point out that the multiplicative representation of a mixed causal-noncausal VAR model does not always exist and implies restrictions on the coefficient matrices $\Phi_{1},...,\Phi_{p}$.
In fact, the set of multiplicative VAR and mixed VAR models are non-nested, i.e. a mixed model does not always admit a multiplicative form and a multiplicative VAR model cannot always be written as a mixed VAR.
@@
with i.i.d. errors, if matrix $\Psi$ is not invertible. Let us now consider the reciprocal problem. We have the following proposition:
Proposition A.1: When $m\geq 2$, a mixed VAR(1) representation with at least one causal and one non-causal roots cannot be factorized as in (a.7).
Proof: For ease of exposition, we consider below the bivariate case $m=2$. The results remain valid for $m\geq 2$. Let us consider a mixed VAR(1) model
$Y_{t}=\Phi_{1}Y_{t-1}+\epsilon_{t},$
where $\Phi_{1}$ has eigenvalues $\lambda_{1}<1$ and $\lambda_{2}>1$. If it can be written in a multiplicative form (a.7), we deduce from the stationarity condition of the process that in this decomposition $det(Id-\Phi z)$ [resp. $det(Id-\Psi z)$] has a unique root $\lambda_{1}$ (resp. $1/\lambda_{2}$).
However, we have: $det(Id-\Phi z)=1-zTr\Phi+z^{2}det\Phi,$ and $det(Id-\Psi z)=1-zTr\Psi+z^{2}det\Psi.$ These polynomials have a unique root if and only of $det\Phi=det\Psi=0$, that is if $\Phi$ and $\Psi$ are non-zero and their rank is equal to 1. Then, we have:
$-\Psi Y_{t}+(Id+\Phi\Psi)Y_{t-1}-\Phi Y_{t-2}=\epsilon_{t}^{*}$
We cannot normalize this equation as in (2.1) since $\Psi$ is not of full rank, and we cannot obtain a VAR(1) model since $\Phi\neq 0$.
Online APPENDICES
B: Predictive Density Estimation
This Appendix describes the kernel-based estimation of the predictive density given in Proposition 1 from the following time series:
$\hat{\varepsilon}_{t}=Y_{t}-\hat{\Phi}_{1}Y_{t-1}-\cdots-\hat{\Phi}_{p}Y_{t-p+1},\;\;t=1,...,T,$
$\hat{Z}_{2,t}=\hat{A}^{2}\left(\begin{array}[]{c}Y_{t}\\ \tilde{Y}_{t-1}\end{array}\right),\;\;t=1,...,T.$
The above time series are used to approximate the density $g$ of $\varepsilon_{t}$ and density $l_{2}$ of $Z_{2,t}$ as follows:
$\hat{g}_{T}(\varepsilon)=\frac{1}{T}\frac{1}{h_{1}^{m}}\sum_{t=1}^{T}K_{m}\left(\frac{\varepsilon-\hat{\varepsilon}_{t}}{h_{1}}\right),$
and
$\hat{l}_{2,T}(z_{2})=\frac{1}{T}\frac{1}{h_{2}^{n_{2}}}\sum_{t=1}^{T}K_{n_{2}}\left(\frac{z_{2}-\hat{Z}_{2,t}}{h_{2}}\right),$
where $h_{1},h_{2}$ are the bandwidths and $K_{m},K_{n_{2}}$ are multivariate kernels of dimensions $m$ and $n_{2}$, respectively. Then, the estimated predictive density is: